# Separating subsets from their images

### v1

@@ 5. $(v)$
We end Section 8.1 on a sour note. An attentive reader might have noticed that we have not discussed case $(l)$ in Table 1, which is the last step to complete the proof of J. In this case, the socle of $G$ is isomorphic to $\Omega^{\epsilon}_{2d}(2^{f})$, and the domain consists of nonsingular $1$-subspaces of the natural module of $G$. As stated in Remark 11, our current method does not give us any control on $1$-dimensional subspaces, and we have not developed an ad hoc approach for this scenario. Hence, the bound that appears in J is actually the general bound from D.
@@ 8.2. Simple diagonal type
Let us first recall the definition of primitive groups of simple diagonal type. Let $T$ be a nonabelian simple group, and let $k\geq 3$ be an integer. Note that we can naturally define three actions on $T^{k}$ – the right-regular permutation representation of $T^{k}$, the component-wise action of $\mathrm{Aut}(T)$, and the action of $\mathrm{Sym}(k)$ permuting indices. All these actions preserve the diagonal of $T^{k}$,
@@
hence we can consider their action on the domain $\Omega=T^{k}/\mathrm{diag}(T^{k})$. For simplicity, we can identify $\Omega$ with $T^{k-1}$ via the set of representatives
@@
Proof of K.
In view of Lemma 12, we can concentrate on $G=T^{k}\rtimes(\mathrm{Out}(T)\times S_{k})$, the largest primitive group of diagonal type with socle $T^{k}$ acting on $T^{k-1}$. We will build a set $A\subseteq T^{k-1}$ that cannot be separated by any of its images under the action of $G$.
@@
$1\in B$, and for every $t\in T$, $B\cap Bt$ is nonempty. (Recall that its existence is guaranteed by E.) Define
@@
It remains to prove that the set $A$ satisfies our conditions, that is, for an arbitrary $g\in G$ we have that $A\cap A^{g}$ is nonempty. For every $g\in G$, there exist $a\in\mathrm{Sym}(k)$, $b\in\mathrm{Out}(t)$, $c\in T$ and $d\in T^{k-1}$ such that $g=abcd$. We claim that, by construction, $A$ is stabilized by $\mathrm{Sym}(k)$, by $\mathrm{Out}(T)$ and by the left semiregular action of $T$. The claim is clear for the last two items, while we shall prove it for the former. Recall that $\mathrm{Sym}(k)$ acts on $T^{k}$ by permuting components, and hence, for every point $(t_{1},\ldots,t_{k-1},1)\mathrm{diag}(T^{k})$ and for every $a\in\mathrm{Sym}(k)$,
@@
(For a better readability, we write the action of $\mathrm{Sym}(k)$ on the indices as a right multiplication. Moreover, to avoid a cumbersome notation, we identify $1=t_{k}$.) Observe that every direct factor of $A$ is equal, and hence the previous formula proves the claim, because
@@
The asymptotic estimate for $\mathbf{F}_{\mathcal{C}}(n)$ are obtained by restricting the attention to the groups of simple diagonal type whose socle is isomorphic to a $k$-fold Cartesian product of groups of Lie type. (Actually, groups of Lie type of characteristic $2$ are those for which the bound on $|\mathrm{Out}(T)|$ cannot be improved, see [25], and thus meet the maximum in $\mathbf{F}_{\mathcal{C}}(n)$.) ∎
@@
Moreover, in the proof of K, our starting choice of $A_{0}$ implies that the size of $A$ depends on $k$ as an exponential function. Meanwhile, if we would have chosen $A_{0}$ to be any difference basis of the socle of appropriate size, by imposing the stability of $A$ under the action of $\mathrm{Sym}(k)$, the constant would have been factorial in $k$. Perhaps, a similar trick could extend our approach beyond the permutation groups of simple diagonal type – but there is no reason not to suspect that the trick would be specific to some limited family of socles, rather than an arbitrary semisimple group.

### v2

@@ 5. $(v)$
We end Section 8.1 on a sour note. An attentive reader might have noticed that we have not discussed case $(l)$ in Table 1, which is the last step to complete the proof of K. In this case, the socle of $G$ is isomorphic to $\Omega^{\epsilon}_{2d}(2^{f})$, and the domain consists of nonsingular $1$-subspaces of the natural module of $G$. As stated in Remark 10, our current method does not give us any control on $1$-dimensional subspaces, and we have not developed an ad hoc approach for this scenario. Hence, the bound that appears in K is actually the general bound from E.
@@ 8.2. Simple diagonal type
Let us first recall the definition of primitive groups of simple diagonal type. Let $T$ be a nonabelian simple group, and let $k\geq 3$ be an integer. Note that we can naturally define three actions on $T^{k}$ – the right-regular permutation representation of $T^{k}$, the component-wise action of $\mathrm{Aut}(T)$, and the action of $\mathrm{Sym}(k)$ permuting indices. We can consider the action induces by these actions on the domain $\Omega=T^{k}/\mathrm{diag}(T^{k})$, where
@@
Observe that the latter two actions preserve the diagonal of $T^{k}$, hence they fall in the stabilser of a point. For simplicity, we can identify $\Omega$ with $T^{k-1}$ via the set of representatives
@@
Proof of L.
In view of Lemma 11, we can concentrate on $G=T^{k}\rtimes(\mathrm{Out}(T)\times S_{k})$, the largest primitive group of diagonal type with socle $T^{k}$ acting on $T^{k-1}$. We will build a set $A\subseteq T^{k-1}$ that is not self-separable.
@@
$1\in B$, and for every $t\in T$, $B\cap Bt$ is nonempty. (Recall that its existence is guaranteed by F.) Define
@@
It remains to prove that the set $A$ satisfies our conditions, that is, for an arbitrary $g\in G$ we have that $A\cap A^{g}$ is nonempty. For every $g\in G$, there exist $a\in\mathrm{Sym}(k)$, $b\in\mathrm{Out}(t)$, $c\in T$ and $d\in T^{k-1}$ such that $g=abcd$. We claim that, by construction, $A$ is stabilized by $\mathrm{Sym}(k)$, by $\mathrm{Out}(T)$ and by the left semiregular action of $T$. The claim is clear for the latter two actions, while we shall prove it for $\mathrm{Sym}(k)$. Recall that $\mathrm{Sym}(k)$ acts on $T^{k}$ by permuting components, and hence, for every point $(t_{1},\ldots,t_{k-1},1)\mathrm{diag}(T^{k})$ and for every $a\in\mathrm{Sym}(k)$,
@@
(For simplicity, we denote the action of $\mathrm{Sym}(k)$ on the indices as right multiplication. Moreover, to avoid a cumbersome notation, we identify $1=t_{k}$.) Observe that every direct factor of $A$ is equal, and hence the previous formula proves the claim, because
@@
The asymptotic estimates for $\mathbf{F}_{\mathcal{C}}(n)$ are obtained by restricting the attention to the groups of simple diagonal type whose socle is isomorphic to a $k$-fold Cartesian product of groups of Lie type. (Actually, groups of Lie type of characteristic $2$ are those for which the bound on $|\mathrm{Out}(T)|$ cannot be improved, see [22], and thus meet the maximum in $\mathbf{F}_{\mathcal{C}}(n)$.) ∎
@@
Moreover, in the proof of L, our starting choice of $A_{0}$ implies that the size of $A$ depends on $k$ as an exponential function. Meanwhile, if we would have chosen $A_{0}$ to be any difference basis of the socle of appropriate size, by imposing the stability of $A$ under the action of $\mathrm{Sym}(k)$, the constant would have been factorial in $k$. Perhaps, a similar trick could extend our approach beyond the permutation groups of simple diagonal type – but there is no reason not to suspect that the trick would be specific to some limited family of socles, rather than an arbitrary semisimple group.